# Integrating Various Software Artifacts for Better LLM-based Bug Localization and Program Repair

### v3

@@ Abstract.
To investigate this and mimic the way human developers fix bugs, we propose DEVLoRe (short for DEVeloper Localization and Repair). In this framework, LLMs first use issue content (description and discussion) and stack error traces to localize buggy methods, then rely on debug information in buggy methods and issue content and stack error to localize buggy lines and generate valid patches. We evaluated the effectiveness of issue content, error stack traces, and debugging information in bug localization and automatic program repair. Our results show that while issue content and error stack is particularly effective in assisting LLMs with fault localization and program repair respectively, different types of software artifacts complement each other in addressing various bugs. By incorporating these three types of artifacts and using the Defects4J v2.0 dataset for evaluation, DEVLoRe successfully localizes 49.3% of single-method bugs and generates 56.0% plausible patches. Additionally, DEVLoRe can localize 47.6% of non-single-method bugs and generates 14.5% plausible patches. Moreover, our framework streamlines the end-to-end process from buggy source code to a complete repair, and achieves a 39.7% and 17.1% of single-method and non-single-method bug repair rate, outperforming current state-of-the-art APR methods. Furthermore, we re-implemented and evaluated our framework, demonstrating its effectiveness in its effectiveness in resolving 9 unique issues compared to other state-of-the-art frameworks using the same or more advanced models on SWE-bench Lite. We also discussed whether a leading framework for Python code can be directly applied to Java code, or vice versa. The source code and experimental results of this work for replication are available at https://github.com/XYZboom/DEVLoRe.
Large Language Model, Automatic Program Repair, Fault Localization, A
@@ 1. Introduction
    We evaluated our framework on SWE-bench Lite and demonstrated DEVLoRe’s effectiveness in resolving 9 unique issues compared to other state-of-the-art frameworks using the same or more advanced models. We further discussed whether a leading framework for Python code can be directly applied to Java code, or vice versa..
@@ 2.1. Debugging Information to the Rescue
The Lang 1b bug involves the createNumber method, which takes a string as its parameter and returns a number represented by that string. While studying the Lang 1b bug, we noticed that the newly added failing test case is the input “0x80000000” for the createNumber method. When this test case is provided, the buggy createNumber method incorrectly treats the input string as an integer, when it should be a long integer instead. The code snippet in Figure 1 shows a portion of the buggy createNumber method in Lang 1b. The local variable hexDigits represents the valid hexadecimal digits. As human software developers, we can quickly recognize that at Line 471, the condition should be “hexDigits > 8 || (hexDigits == 8 && firstSigDigit > ’7’)” (firstSigDigit refers to the first valid hexadecimal digit).
@@
For this bug, we provided the LLM with Lang 1b’s issue content (https://issues.apache.org/jira/browse/LANG-747) and the stack error message shown in Figure 2, but both could not help the LLM generate a plausible fix or near-to-correct patch. This is because the leap required in reasoning is significant, and the buggy method is quite long. Neither of the information provided above could pinpoint the buggy line precisely, let alone generate a plausible fix. To address this, we introduced debugging information to help the LLM understand the necessary changes. One piece of debug information we provided was a series of variable-value pairs, commons.lang3.math.NumberUtils:createNumber:468 {hexDigits:8}. This indicates that the code is about to execute Line 468 and that the value of hexDigits is 8 for the given test case. This enables the LLM to understand why the code on Line 474 was executed instead of Line 471. In one of the responses, the LLM suggested changing Line 468 to if (hexDigits > 16 || (hexDigits == 8 && str.charAt(2) >= ’8’)) , which, although not completely accurate, demonstrates that the provided debug information has a positive impact on fixing the bug.
@@ 2.2. Issue Content to the Rescue
The debug information, org.apache.commons.lang.StringUtils:containsIgnoreCase:1045 {“str”:“ß”, “searchStr”:“SS”}, and the error stack trace shown in Figure 3 merely re-display the test cases and do not clearly highlight the relationship between the character 0x00DF and ’SS’. As a result, neither the error stack trace nor the debug information provides any useful clues about the bug, and LLMs cannot generate a plausible fix when provided with either the stack message or the debug information alone. However, with the description and expected behavior outlined in the issue content, GPT-4o is able to generate a correct patch.
@@ 5.1.2. Localizing Buggy Line(s)
To evaluate LLMs’ ability in localizing buggy lines, we provide the LLM with the buggy method(s) baseline (i.e., the ground truth which can be extracted from human developer’s correct patch) and ask the LLM to predict which lines are buggy. If the line number generated by the LLM exactly matches the first line added in the human developer’s correct patch in Defects4J, we consider it an exact match. If the predicted line number falls within a range of $n$ lines from the first added line in the human developer’s patch (for example, if the correct patch starts at Line 368, and the LLM’s prediction falls within the range from $368-n$ to $368+n$), we classify it as a Range-n match.
@@
[…]
@@ Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.

### v4

@@ Abstract.
To investigate this and mimic the way human developers fix bugs, we propose DEVLoRe (short for DEVeloper Localization and Repair). In this framework, LLMs first use issue content (description and discussion) and stack error traces to localize buggy methods, then rely on debug information in buggy methods and issue content and stack error to localize buggy lines and generate valid patches. We evaluated the effectiveness of issue content, error stack traces, and debugging information in bug localization and automatic program repair. Our results show that while issue content and error stack is particularly effective in assisting LLMs with fault localization and program repair respectively, different types of software artifacts complement each other in addressing various bugs. By incorporating these three types of artifacts and using the Defects4J v2.0 dataset for evaluation, DEVLoRe successfully localizes 49.3% of single-method bugs and generates 56.0% plausible patches. Additionally, DEVLoRe can localize 47.6% of non-single-method bugs and generates 14.5% plausible patches. Moreover, our framework streamlines the end-to-end process from buggy source code to a complete repair, and achieves a 39.7% and 17.1% of single-method and non-single-method bug repair rate, outperforming current state-of-the-art APR methods. Furthermore, we re-implemented and evaluated our framework, demonstrating its effectiveness in resolving 9 unique issues compared to other state-of-the-art frameworks using the same or more advanced models on SWE-bench Lite. We also discussed whether a leading framework for Python code can be directly applied to Java code, or vice versa. The source code and experimental results of this work for replication are available at https://github.com/XYZboom/DEVLoRe.
Large Language Model, Automatic Program Repair, Fault Localization
@@ 1. Introduction
    We evaluated our framework on SWE-bench Lite and demonstrated DEVLoRe’s effectiveness in resolving 9 unique issues compared to other state-of-the-art frameworks using the same or more advanced models. We further discussed whether a leading framework for Python code can be directly applied to Java code, or vice versa.
@@ 2.1. Debugging Information to the Rescue
The Lang 1b bug involves the createNumber method, which takes a string as its parameter and returns a number represented by that string. While studying the Lang 1b bug, we noticed that the newly added failing test case is the input “0x80000000” for the createNumber method. When this test case is provided, the buggy createNumber method incorrectly treats the input string as an integer, when it should be a long integer instead. The code snippet in Figure 1 shows a portion of the buggy createNumber method in Lang 1b. The local variable hexDigits represents the valid hexadecimal digits. As human software developers, we can quickly recognize that at Line 471, the condition should be “hexDigits > 8 || (hexDigits == 8 && firstSigDigit > ‘7’)” (firstSigDigit refers to the first valid hexadecimal digit).
@@
For this bug, we provided the LLM with Lang 1b’s issue content (https://issues.apache.org/jira/browse/LANG-747) and the stack error message shown in Figure 2, but both could not help the LLM generate a plausible fix or near-to-correct patch. This is because the leap required in reasoning is significant, and the buggy method is quite long. Neither of the information provided above could pinpoint the buggy line precisely, let alone generate a plausible fix. To address this, we introduced debugging information to help the LLM understand the necessary changes. One piece of debug information we provided was a series of variable-value pairs, commons.lang3.math.NumberUtils:createNumber:468 {hexDigits:8}. This indicates that the code is about to execute Line 468 and that the value of hexDigits is 8 for the given test case. This enables the LLM to understand why the code on Line 474 was executed instead of Line 471. In one of the responses, the LLM suggested changing Line 468 to if (hexDigits > 16 || (hexDigits == 8 && str.charAt(2) >= ‘8’)) , which, although not completely accurate, demonstrates that the provided debug information has a positive impact on fixing the bug.
@@ 2.2. Issue Content to the Rescue
The debug information, org.apache.commons.lang.StringUtils:containsIgnoreCase:1045 {“str”:“ß”, “searchStr”:“SS”}, and the error stack trace shown in Figure 3 merely re-display the test cases and do not clearly highlight the relationship between the character 0x00DF and ‘SS’. As a result, neither the error stack trace nor the debug information provides any useful clues about the bug, and LLMs cannot generate a plausible fix when provided with either the stack message or the debug information alone. However, with the description and expected behavior outlined in the issue content, GPT-4o is able to generate a correct patch.
@@ 4.2. Parameters in Experiments and Baseline Selection
4.2.1. Parameters in Experiments:
Due to the extremely long debugging information for some bugs recorded by DebugRecorder, along with issue content and error stack, which exceeds the token limit of most LLMs, we opted for the GPT-4 series model, known for its ability to handle long token sequences with a 128k token context window. To manage costs, we selected the cost-effective GPT-4o-mini model, also with a 128k token context window, and used the default settings (Temperature=0.5, Top-p=1). For the experimental setup, before automating the entire process, we manually used a client with access to multiple LLMs and provided artifact information for 10 random selected Defects4J bugs to evaluate the feasibility of our approach and to select appropriate models—considering factors such as low cost, a large token context window, and minimal hallucination rate. During this step, we found that the GPT-3.5 series exhibited a high rate of hallucination when generating patches, and we excluded it from consideration. We also determined the appropriate format for the artifacts to be fed into the LLM. For example, we pruned the debug information such as only recording variables that have changed, cropping long arrays and lists to minimize token size. Additionally, we instructed the LLM to generate *SEARCH/REPLACE* edits when producing patches, in order to reduce hallucination rate, as adopted by SOTA approaches (Xia et al., 2024a; Ouyang et al., 2024; Zhang et al., 2024; Liu et al., 2024). After selecting the model, prompts and artifacts format, we used this preliminary process to determine the optimal number of LLM calls at each step: localizing buggy methods, localizing buggy lines, and generating patches. The rationale is that increasing the number of LLM calls beyond a certain point does not lead to significant performance gains. Based on these manual experiments, we finalized the settings of the parameters. For buggy method localization, we call the LLM only once. For buggy line localization, we call the LLM 10 times, collecting 10 responses, each identifying potential buggy lines. We then filter out duplicate responses (i.e., identical buggy line locations). Finally, for each unique response, we ask the LLM to generate patches three times, ensuring that multiple potential fixes are considered.
4.2.2. Baseline Selections:
We selected state-of-the-art LLM-based fault localization and program repair frameworks on the Defects4J dataset as our baselines for two reasons. First, using the same Defects4J dataset allows us to directly compare our results with those reported in the baseline papers. Second, we focus on LLM-based approaches to better evaluate our framework, which is also LLM-based and leveraging a set of software artifacts.
4.3. Research Questions
In this work, we aim to answer the following research questions (RQs) for evaluating the effectiveness of various software artifacts in assisting the chosen LLM in bug localization and APR task.
  • RQ1. Which software artifacts can better assist LLMs in localizing buggy methods and buggy lines when provided with source code? This RQ aims to identify the specific types of software artifacts - issue content, debugging information and error stack trace - that enhance the ability of LLMs to accurately pinpoint methods and code that contain bugs.
  • RQ2. Which software artifacts can better assist LLMs in generating plausible patches when provided with buggy methods? By answering this RQ, we can identify the specific types of software artifacts in assisting LLMs to produce plausible patches which can pass all unit tests.
  • RQ3. What is the overall performance of DEVLoRe in bug localization and program repair? Unlike some LLM-based approaches that focus on either fault localization or program repair while relying on traditional methods such as spectrum-based approaches (Abreu et al., 2007) for fault localization (Li et al., 2025; Xia et al., 2023a), DEVLoRe supports an end-to-end bug identification and program repair workflow. This RQ aims to evaluate the overall performance of DEVLoRe from the initial input of a code repository to the final output of a complete repair.
@@ 5.1.2. Localizing Buggy Line(s)
To evaluate LLMs’ ability in localizing buggy lines, we provide the LLM with the buggy method(s) baseline (i.e., the ground truth which can be extracted from human developer’s correct patch) and ask the LLM to predict which lines are buggy. If the line number generated by the LLM exactly matches the first line added in the human developer’s correct patch in Defects4J, we consider it an exact match. If the predicted line number falls within a range of $n$ lines from the first added line in the human developer’s patch (for example, if the correct patch starts at Line 368, and the LLM’s prediction falls within the range from 368n368-n to $368+n$), we classify it as a Range-n match.
@@
Evaluate Plausible Patches by the Defects4J framework $\approx{}$300s
Total $\approx{}$315s
@@ Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311 and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.